PDCL1MAA—Macros for using Latin-1 characters with Malvern A

Version 1.3, last changed pdc 1995–03–30

P. Damian Cugley

1. Introduction

This document describes pdcl1maa.tex, a collection of definitions designed to make Malvern A characters available using the ISO 8859–1 input encoding (Latin-1).

1.1. About this document The definitions file and the printed documentation are both in pdcl1maa.dtx, a ‘documented TEX macros’ file* which, when processed by plain TEX, generates a fresh copy of pdcl1maa.tex in the current directory in addition to the dvi file. This ensures that the printed documentation and the code it describes are identical. The dtx file uses the mac-

ros in pdccode.tex.

2. File identification

Nowadays, macro files start with some comments identifying the file.

 1 % pdcl1maa.tex 1.3 1995/03/30 -- Latin-1 chars in Malvern-A

2

 3 %%%@TeX-definition-file {
 4 %%% filename       = "$texmf/tex/plain/pdcmac/pdcl1maa.tex",
 5 %%% version        = "1.3",
 6 %%% date           = "1995/03/30",
 7 %%% package        = "pdcmac 1.0",
 8 %%% author         = "P. Damian Cugley",
 9 %%% email          = "damian.cugley@comlab.ox.ac.uk",
10 %%% address        = "Oxford University Computing Laboratory,
11 %%%                   Parks Road, Oxford  OX1 3QD, UK",
12 %%% codetable      = "USASCII",
13 %%% keywords       = "TeX, plain TeX, macros",
14 %%% supported      = "Maybe",
15 %%% abstract       = "ISO 8859-1 characters using Malvern-A fonts.
16 %%%                   This file was generated by running
17 %%%                   plain TeX on pdcl1maa.dtx",
18 %%% copyright      = "Copyright (c) 1995 P. Damian Cugley",
19 %%% copying        = "DO NOT DISTRIBUTE THIS FILE.
20 %%%                   Distribute pdcl1maa.dtx only as part of the
21 %%%                   package it came in.",
22 %%% dependencies   = "",
23 %%% }

24

25 \message{1.3 <pdc 1995/03/30>}

26

3. ISO 8859–1 input

3.1. Symbols Those characters whose code is the same as the code of the corresponding glyph are left alone (with catcode 12). For example, ^^a3 (pounds sterling) is left alone.

* The LaTEX 2e distribution uses files with the ‘dtx’ suffix for similar purposes.

27–71 PDCL1MAA—Macros for using Latin-1 characters with Malvern˜ A 1.3 1995–03–30

Those that have an equivalent \chardef (or \mathchardef) name are defined with \let. Those with an equivalent standard ligature use that ligature.

I don’t use \declareactivechar on these because they would not work properly in verbatim text with catcode 12. (Extending verbatim text to include Latin-1 will take a few definitions along the same lines as \ttlq etc.)

27 \catcode` ^^a0\active \let^^a0=~
28 \catcode` ^^a5\active \let^^a5=\yen
29 \catcode` ^^a7\active \let^^a7=\S
30 \catcode` ^^a8\active \let^^a8=\"
31 \catcode` ^^a9\active \let^^a9=\copyright
32 \catcode` ^^aa\active \let^^aa=\orda
33 \catcode` ^^ab\active \def^^ab{<<}
34 \catcode` ^^ac\active \let^^ac=\lnot
35 \catcode` ^^ad\active \let^^ad=\-
36 \catcode` ^^ae\active \let^^ae=\registered
37 \catcode` ^^af\active \let^^af=\=
38 \catcode` ^^b0\active \let^^b0=\degrees
39 \catcode` ^^b1\active \let^^b1=\pm
40 \catcode` ^^b2\active \def^^b2{^2}
41 \catcode` ^^b3\active \def^^b3{^3}
42 \catcode` ^^b4\active \let^^b4=\´
43 \catcode` ^^b5\active \let^^b5=\mu
44 \catcode` ^^b6\active \let^^b6=\P
45 \catcode` ^^b7\active \let^^b7=\gbdecimal
46 \catcode` ^^b8\active \let^^b8=\c
47 \catcode` ^^b9\active \def^^b9{^1}
48 \catcode` ^^ba\active \let^^ba=\ordo
49 \catcode` ^^bb\active \def^^bb{>>}
50 \catcode` ^^bc\active \def^^bc{\frac14}
51 \catcode` ^^bd\active \def^^bd{\frac12}
52 \catcode` ^^be\active \def^^be{\frac34}
53 \catcode` ^^bf\active \def^^bf{?` }
54 \catcode` ^^d7\active \let^^d7=\times
55 \catcode` ^^f7\active \let^^f7=\div

3.2. Composite letters

The characters that represent marks for composite letters* are \let to the corresponding composite-letter-making command—one may write ‘¨ u¸ c’ to get ‘üç’.

56 \catcode` ^^c0\active \def^^c0{\` A}
57 \catcode` ^^c1\active \def^^c1{\´ A}
58 \catcode` ^^c2\active \def^^c2{\^A}
59 \catcode` ^^c3\active \def^^c3{\~A}
60 \catcode` ^^c4\active \def^^c4{\"A}
61 \catcode` ^^c5\active \let^^c5=\AA
62 \catcode` ^^c6\active \let^^c6=\AE
63 \catcode` ^^c7\active \def^^c7{\c C}
64 \catcode` ^^c8\active \def^^c8{\` E}
65 \catcode` ^^c9\active \def^^c9{\´ E}
66 \catcode` ^^ca\active \def^^ca{\^E}
67 \catcode` ^^cb\active \def^^cb{\"E}
68 \catcode` ^^cc\active \def^^cc{\` I}
69 \catcode` ^^cd\active \def^^cd{\´ I}
70 \catcode` ^^ce\active \def^^ce{\^I}
71 \catcode` ^^cf\active \def^^cf{\"I}

* This phrase is used on the grounds that the word accent tends to provoke lectures on ways in which many composite letters do not represent accents in the languages in which they are used.

72–117 PDCL1MAA—Macros for using Latin-1 characters with Malvern A 1.3 1995–03–30

72 \catcode` ^^d0\active \let^^d0=\DH
73 \catcode` ^^d1\active \def^^d1{\~N}
74 \catcode` ^^d2\active \def^^d2{\` O}
75 \catcode` ^^d3\active \def^^d3{\´ O}
76 \catcode` ^^d4\active \def^^d4{\^O}
77 \catcode` ^^d5\active \def^^d5{\~O}
78 \catcode` ^^d6\active \def^^d6{\"O}
79 \catcode` ^^d8\active \let^^d8=\O
80 \catcode` ^^d9\active \def^^d9{\` U}
81 \catcode` ^^da\active \def^^da{\´ U}
82 \catcode` ^^db\active \def^^db{\^U}
83 \catcode` ^^dc\active \def^^dc{\"U}
84 \catcode` ^^dd\active \def^^dd{\´ Y}
85 \catcode` ^^de\active \let^^de=\TH
86 \catcode` ^^df\active \let^^df=\ss
87 \catcode` ^^e0\active \def^^e0{\´ a}
88 \catcode` ^^e1\active \def^^e1{\` a}
89 \catcode` ^^e2\active \def^^e2{\^a}
90 \catcode` ^^e3\active \def^^e3{\~a}
91 \catcode` ^^e4\active \def^^e4{\"a}
92 \catcode` ^^e5\active \let^^e5=\aa
93 \catcode` ^^e6\active \let^^e6=\ae
94 \catcode` ^^e7\active \def^^e7{\c c}
95 \catcode` ^^e8\active \def^^e8{\` e}
96 \catcode` ^^e9\active \def^^e9{\´ e}
97 \catcode` ^^ea\active \def^^ea{\^e}
98 \catcode` ^^eb\active \def^^eb{\"e}
99 \catcode` ^^ec\active \def^^ec{\` i}
100 \catcode` ^^ed\active \def^^ed{\´ i}
101 \catcode` ^^ee\active \def^^ee{\^i}
102 \catcode` ^^ef\active \def^^ef{\"i}
103 \catcode` ^^f0\active \let^^f0=\dh
104 \catcode` ^^f1\active \def^^f1{\~n}
105 \catcode` ^^f2\active \def^^f2{\` o}
106 \catcode` ^^f3\active \def^^f3{\´ o}
107 \catcode` ^^f4\active \def^^f4{\^o}
108 \catcode` ^^f5\active \def^^f5{\~o}
109 \catcode` ^^f6\active \def^^f6{\"o}
110 \catcode` ^^f8\active \let^^f8=\o
111 \catcode` ^^f9\active \def^^f9{\` u}
112 \catcode` ^^fa\active \def^^fa{\´ u}
113 \catcode` ^^fb\active \def^^fb{\^u}
114 \catcode` ^^fc\active \def^^fc{\"u}
115 \catcode` ^^fd\active \def^^fd{\´ y}
116 \catcode` ^^fe\active \let^^fe=\th
117 \catcode` ^^ff\active \def^^ff{\"y}